# Architecting mechanisms of damage in topological metamaterials

### v2

@@ 4.1 Dominant Lattice behaviour
Figure 4: Dominant behaviour of lattices. (a) Axial energy portion in the entire lattice preceding the first failure event. Visualisation of the affine strain producing mechanism (GH mode) for: (b) RK; (c) LP1; and, (d) LUP1. Comparison of the internal angle of a unit cell from the simulations ($\alpha_{\mathrm{s}}$) with that required by the GH mode ($\alpha_{\mathrm{GH}}$) to displace the boundary the equivalent distance (i.e. $\alpha_{\mathrm{d}}=\alpha_{\mathrm{s}}-\alpha_{\mathrm{GH}}$) for: (e, h) LTK, (f, i) LP1; and, (g, j) RP1 at SR = 3 and 30 respectively. The difference in angle is visualised through the colour directly to the right of each unit cell.
@@
In practice, bonds are constrained from freely rotating around each site due to the local bending resistance associated with finite thickness joints. However, contrasting the idealised GH modes (i.e., affine strain-producing mechanisms) with the lattice deformation in the simulator, offers insights into the mechanics while traversing SR-space. Geometries that would ideally possess a GH mode will, in reality, become bending-dominated at high SR values, if the applied boundary displacements closely match those required by the GH mode (e.g., Fig. 4c). The LTK, LP, and RP lattices, all belong in this category. On the contrary, lattices that either do not posses a GH mode, such as the RK lattice, or whose GH mode does not produce boundary displacements compatible with the simulated ones, such as LUP (Fig. 4d), remain stretching-dominated throughout the SR range. Here, there is a caveat for the RK lattice, whose uniform FM depicted in Fig. 4b is not classified as a GH mode, as it does not produce any affine strain to linear order in the displacements. This mode is rather an infinitesimal increment of the GH mode (Hutchinson and Fleck, 2006), which can only map to boundary displacements in compression through an instability, leading to stretching-dominated behaviour in tension, across the SR range.
@@
To explore whether non-idealised lattices deform along the GH mode, further visual comparisons are presented for low (Fig. 4e-g) and high (Fig. 4h-j) SR. These figures highlight the difference between the simulated local angle, $\alpha_{\mathrm{s}}$, and the expected GH angle, $\alpha_{\mathrm{GH}}$. Despite the majority of the bulk following the GH mode for the bending-dominated lattices at $\text{SR}=30$, they also posses regions that diverge from this behaviour. Bending stresses mirror this trend (Fig. 5a,c,l), with higher stress values appearing in the same regions as the maps for $\alpha_{\mathrm{d}}$ in Fig. 4. Below, we argue that the form of the regions that diverge from the GH mode can be connected to the edge actuation through the bulk-boundary correspondence. Thus, the phenomena observed can be interpreted in light of the compliance added to the system by the pre-crack.
@@
Figure 5: Mapping of bending stresses through the bulk-boundary correspondence. Maximum normalised bending stress in the elements cross section $\sigma_{\mathrm{\textbf{Bend}}}=|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$ and visualisation of FM through the magnitude of the total modulation ($|\phi|$) of the wave spectrum for: (a, e) RK (0.1,0.1,0.1); (b, f) LP2 (-0.02,0.18,0.18); (c, g) LP1 (-0.1,0.1,0.1); (d, h) LP3 (-0.16,0.04,0.04); (l, m) RP1 (0.1,-0.1,-0.1); and, (n, o) RP2 (0.02,-0.18,-0.18). (i) Normalised bending stress along the arc length. The bending stress is normalised with the peak bending stress along this length. The arc length is illustrated in (a). (j) Decay spectrum of zero modes on a horizontal surface; and, (k) associated wave angles. All figures associated with simulations are for SR = 30.
@@
The bending stresses along the mid-plane (Fig. 5i) reveal the FM actuation both at the crack edge and within the lattice. For LTK and LP1, the actuation is mainly concentrated in a sharp peak at the crack tip, while for RP1, the peak remains close to the uniform value observed in the bulk. All lattices are relaxed along the rest of the crack edge, to the left of the crack tip. By modeling the FM as a modulation of the uniform zero-energy GH mode (Rocklin et al., 2017), without restricting the description to linear order, we can leverage bulk-boundary correspondence, prescribing an analogue of the observed dominant actuation pattern along the mid-plane, to visualize the resulting FM amplitude, $|\phi|$, in the bulk. In particular, we prescribe a Gaussian spike of unit height and width, centred at the crack-tip, for the LP/LTK lattices (Fig. 5e-h), and a sigmoid function centred at the crack-tip for RP (Fig. 5m,o). Different prescriptions can be superimposed, so one could easily add the Gaussian and sigmoid solutions together with different relative amplitudes, to reach a more accurate visualisations of the FM in the bulk for each lattice. However, the goal here is to capture the most dominant components of the disturbance; the method should be seen as a simple heuristic (further details of the visualisation can be found in B).
@@
Three additional lattices within the LP/RP ($\mp$,$\pm$,$\pm$) equivalence classes are simulated to further illustrate how the form of the FM influence the stress profile within the lattice. Even with the simplified aforementioned prescriptions, the FM fields in the bulk (Fig. 5e-h,m,o) map qualitatively well onto the major components of the bending stresses in the lattice (Fig. 5a-d,l,n). To better understand the form of the FM in the bulk, Fig. 5j-k, can offer valuable insights. Firstly, the penetration into the bulk of the localised disturbance can be connected to the decay rate $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})$ spectrum. At the homogenisation limit (i.e small $q_{\mathrm{1}}$), non-polarised lattices (e.g LTK) are expected to decay faster because they have $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})\sim\mathcal{O}(q_{\mathrm{1}})$, compared to $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})\sim\mathcal{O}(q^{2}_{\mathrm{1}})$ for polarised lattices (Kane and Lubensky, 2014). Here, we observe that, in general, the magnitude of $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})$ across the whole range is more indicative, which makes sense for a localised disturbance, as it samples a wide spectrum in $q_{\mathrm{1}}$-space. Subsequently, LTK and LP2 decay quickly into the bulk, with the deformation away from the crack-tip reverting back to that associated with the idealised GH mode. Contrarily, the spectra of LP1 and LP3 exhibit lower decay rates, with the disturbances influencing the behaviour deeper into the bulk.
@@
As the large $q_{\mathrm{1}}$ components of the mode decay faster, the direction of propagation for the disturbance far from the crack tip is dictated by the small $q_{\mathrm{1}}$ limit of $\theta(q_{\mathrm{1}})$. This is clearly discernible by comparing, e.g., LP2 and LP3, where the propagation for the former happens at an angle much closer to the horizontal. Additionally, the form of the extra modes for RP, which propagate towards the left edge (positive $\theta$), do not offer an alternative deformation mechanism for the units within the bulk. The only mode available for these is associated with the uniform GH mode, leading to no significant stress localization at the crack tip. However, notable reduction in bending stress is observed along the right edge of RP1-2 (Fig. 5l,n), due to that edge’s FM which locally increase the compliance.
@@
Finally, the stress within the bulk of stretching-dominated lattices (e.g., LUP1) is not significantly affected by the FM introduced at the crack edge. However, at the crack-tip the FM are actuated in a localised manner (Fig. 5i) and these greatly influence the failure, which will be discussed in the following section.
4.2 Dominant Lattice Failure
@@
Fig. 6a provides the axial stress fraction in the element that fails in the step immediately before fracture. The damage paths, provided in Fig. 6b-g. each showing results for a range of SR, appear to be influenced by the proportion of axial and bending stress within the element. When bending or stretching dominates the fracture, the crack appears to propagate horizontally parallel to the pre-crack edge for symmetric about the mid-plane lattices, i.e. RK at SR = 6, LTK at SR = 2-30, LP1 at SR = 6-30, and RP1 at SR = 20-30. As the stress contributions become similar, the damage path generally deviates from the pre-crack plane as it traverses the lattice.
@@
Figure 6: Dominant failure of lattices. (a) Axial stress fraction in member that fractures preceding the first failure event. (b) The damage process is visualised with a line traversing the domain, connecting separated nodes (fracture locations) within the lattice, indicated by yellow markers. For clarity we do not show the microstructure throughout the domain. The damage process for SR = 3, 6, 10, 30 are plotted with green, purple, black, and blue lines and associated rectangular markers respectively (refer to (a)). Arrows indicate the direction of damage and circular markers indicate a discontinuity in the damage path. Damage paths for: (c) RK; (d) LTK; (e) LP1; (f) RP1; and, (g) LUP1.
@@
While the LUP1 global behaviour is stretching-dominated, the fracture of its elements is strongly influenced by the bending (Fig. 6a) associated with the FM localising at the pre-crack edge at large SR (Fig. 7a,d). As such, strong alignment of the damage path with the long-wavelength limit of $\theta_{\mathrm{2}}$ (Fig. 7i) is observed. Simulating other lattices from the equivalence class (-,+,-) at SR = 30, it indicates this is a robust phenomenon: for LUP2 (-0.18,0.04,-0.1) and LUP3 (-0.16,0.04,-0.16) the damage path (Fig. 7g) follows the long-wavelength limit of $\theta_{2}$ and $\theta_{1}$ respectively (Fig. 7d-f,i). This results in the damage evolution following a shallower angle than LUP1. Close to the long-wavelength limit, $\theta_{1}$ propagates further into the bulk for LUP1-3. While the lattice appears to favour propagating the crack in the direction with the smallest $\kappa^{\mathrm{FM}}$ (i.e. LUP3), if $\theta$ is behind the crack tip, it instead favours the modes that propagate forward, ahead of the crack tip (i.e. LUP1 and 2). This phenomenon is also observed for the RDP lattices, provided in the C. For lattices in the ($\pm$$\mp$,$\pm$) equivalence classes, at large SR, there is a strong alignment of the damage path with the FM localised at the crack tip, providing a control parameter to steer the localised fracture of its underlying microstructure. As the SR reduces, the crack tends towards a horizontal path (Fig. 6g), with these FM having a reduced influence on the damage.
@@
Figure 7: Damage path control using FM. Maximum normalised bending stress in the elements cross section $\sigma_{\mathrm{\textbf{Bend}}}=|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$ and visualisation of FM through the magnitude of the total modulation ($|\phi|$) of the wave spectrum for: (a, d) LUP1 (-0.1,0.1,-0.1); (b, e) LUP2 (-0.18,0.04,-0.1); and, (c, f) LUP3 (-0.16,0.04,-0.16). (g) Damage path for LUP1, LUP2, LUP3 at SR = 30. (h) Damage path for lattice with a domain wall separating RDP1 and LUP1. (i) $\theta$ spectrum for LUP1, LUP2 and LUP3. All figures associated with simulations are for SR = 30.
5 Effect of Discontinuities
@@
5.1 Topological domain walls
@@
Figure 8: Effect of domain walls. Normalised applied stress vs nominal strain for: (a) SR = 3; and, (b) SR = 30. Maximum normalised stress in the elements cross section for D-FM, D-SSS1 and D-SSS2 at specified nominal strains for: (i,iii) SR = 3; and, (ii,iv-viii) SR = 30. Where $\sigma_{\mathrm{M}}=\sigma_{\mathrm{A}}/\sigma_{\mathrm{A}}^{\mathrm{u}}\pm|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$. In (v), (vi), and (viii) yellow stars indicated nodes that are added into the system due to the separation of elements (element fracture) along the domain wall. For clarity, an insert is also included where circular yellow markers indicate these nodes.
Pairing lattices from different equivalence classes provides a mechanism to arrest the crack at the domain wall. At SR = 3, LP1 has a higher peak stress than RP1 (Fig. 8b). Once initiated, the crack traverses D-FM until it meets the domain wall (Fig. 8i). The load then increases before the damage continues, resulting in the stiffening behaviour observed in Fig. 8a. In contrast, at SR = 30, RP1 has a higher peak stress, resulting in the damage propagating unstably through D-FM with a sudden reduction in stress (Fig. 8a). This stiffening is also observed for D-SSS1 and D-SSS2 once the damage reaches the SSS domain wall.

### v3

@@ IV.1 Dominant Lattice behaviour
Figure 4: Dominant behaviour of lattices. (a) Axial energy portion in the entire lattice preceding the first failure event. Visualisation of the affine strain producing mechanism (GH mode) for: (b) RK; (c) LP1; and, (d) LUP1. Comparison of the internal angle of a unit cell from the simulations ($\alpha_{\mathrm{s}}$) with that required by the GH mode ($\alpha_{\mathrm{GH}}$) to displace the boundary the equivalent distance (i.e. $\alpha_{\mathrm{d}}=\alpha_{\mathrm{s}}-\alpha_{\mathrm{GH}}$) for: (e, h) LTK, (f, i) LP1; and, (g, j) RP1 at SR = 3 and 30 respectively. The difference in angle is visualised through the colour directly to the right of each unit cell. The simulations for these lattices are included in movie S1.
@@
In practice, bonds are constrained from freely rotating around each site due to the local bending resistance associated with finite thickness joints. However, contrasting the idealised GH modes (i.e., affine strain-producing mechanisms) with the lattice deformation in the simulator, offers insights into the mechanics while traversing SR-space. Geometries that exhibit a GH mode will, in practice, become bending-dominated at high SR values when the applied boundary displacements closely correspond to those demanded by the GH mode (e.g., Fig. 4c). The LTK, LP, and RP lattices, all belong in this category. Conversely, lattices without a GH mode, such as RK, or whose GH mode yields boundary displacements incompatible with the simulations, such as LUP (Fig. 4d), remain stretching-dominated across the entire SR range. Here, there is a caveat for the RK lattice, whose uniform FM depicted in Fig. 4b is not classified as a GH mode, as it does not produce any affine strain to linear order in the displacements. This mode is rather an infinitesimal increment of the GH mode [31], which can only map to boundary displacements in compression through an instability, leading to stretching-dominated behaviour in tension across the SR range.
@@
To explore whether non-idealised lattices deform along the GH mode, further visual comparisons are presented for low (Fig. 4e-g) and high (Fig. 4h-j) SR. These figures highlight the difference between the simulated local angle, $\alpha_{\mathrm{s}}$, and the expected GH angle, $\alpha_{\mathrm{GH}}$. Despite the majority of the bulk following the GH mode for the bending-dominated lattices at $\text{SR}=30$, they also posses regions that diverge from this behaviour. Bending stresses mirror this trend (Fig. 5a,c,i), with higher stress values appearing in the same regions as the maps for $\alpha_{\mathrm{d}}$ in Fig. 4. Below, we argue that the form of the regions that diverge from the GH mode can be connected to the edge actuation through the bulk-boundary correspondence. Thus, the phenomena observed can be interpreted in light of the compliance added to the system by the pre-crack.
@@
Figure 5: Mapping of bending stresses through the bulk-boundary correspondence. Maximum normalised bending stress in the elements cross section $\sigma_{\mathrm{\textbf{Bend}}}=|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$ and visualisation of FM through the magnitude of the total modulation ($|\phi|$) of the wave spectrum for: (a, e) RK (0.1,0.1,0.1); (b, f) LP2 (-0.02,0.18,0.18); (c, g) LP1 (-0.1,0.1,0.1); (d, h) LP3 (-0.16,0.04,0.04); (i, l) RP1 (0.1,-0.1,-0.1); and, (j, m) RP2 (0.02,-0.18,-0.18). (k) Normalised bending stress along the arc length. The bending stress is normalised with the peak bending stress along this length. The arc length is illustrated in (a). (n) Decay spectrum of zero modes on a horizontal surface; and, (o) associated wave angles. All figures associated with simulations are for SR = 30. The simulations for these lattices are included in movie S1-2.
@@
The bending stresses along the mid-plane (Fig. 5k) reveal the FM actuation both at the crack edge and within the lattice. For LTK and LP1, the actuation is mainly concentrated in a sharp peak at the crack tip, while for RP1, the peak remains close to the uniform value observed in the bulk. All lattices are relaxed along the rest of the crack edge, to the left of the crack tip. By modelling the FM as a modulation of the uniform zero-energy GH mode [38], without restricting the description to linear order, we can leverage bulk-boundary correspondence, prescribing an analogue of the observed dominant actuation pattern along the mid-plane, to visualize the resulting FM amplitude, $|\phi|$, in the bulk. In particular, we prescribe a Gaussian spike of unit height and width, centred at the crack-tip, for the LP/LTK lattices (Fig. 5e-h), and a sigmoid function centred at the crack-tip for RP (Fig. 5l,m). Different prescriptions can be superimposed, so one could easily add the Gaussian and sigmoid solutions together with different relative amplitudes, to reach a more accurate visualisations of the FM in the bulk for each lattice. However, the goal here is to capture the most dominant components of the disturbance; the method should be seen as a simple heuristic (further details of the visualisation can be found in Appendix D).
@@
Three additional lattices within the LP/RP ($\mp$$\pm,$$\pm$) equivalence classes are simulated to further illustrate how the form of the FM influence the stress profile within the lattice. Even with the simplified aforementioned prescriptions, the FM fields in the bulk (Fig. 5e-h,l,m) map qualitatively well onto the major components of the bending stresses in the lattice (Fig. 5a-d,i,j). To better understand the form of the FM in the bulk, Fig. 5n,o, can offer valuable insights. Firstly, the penetration into the bulk of the localised disturbance can be connected to the decay rate $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})$ spectrum. At the homogenisation limit (i.e small $q_{\mathrm{1}}$), non-polarised lattices (e.g LTK) are expected to decay faster because they have $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})\sim\mathcal{O}(q_{\mathrm{1}})$, compared to $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})\sim\mathcal{O}(q^{2}_{\mathrm{1}})$ for polarised lattices [20]. Here, we observe that, in general, the magnitude of $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})$ across the whole range is more indicative, which makes sense for a localised disturbance, as it samples a wide spectrum in $q_{\mathrm{1}}$-space. Subsequently, LTK and LP2 decay quickly into the bulk, with the deformation away from the crack-tip reverting back to that associated with the idealised GH mode. Contrarily, the spectra of LP1 and LP3 exhibit lower decay rates, with the disturbances influencing the behaviour deeper into the bulk.
@@
As the large $q_{\mathrm{1}}$ components of the mode decay faster, the direction of propagation for the disturbance far from the crack tip is dictated by the small $q_{\mathrm{1}}$ limit of $\theta(q_{\mathrm{1}})$. This is clearly discernible by comparing, e.g., LP2 and LP3, where the propagation for the former happens at an angle much closer to the horizontal. Additionally, the form of the extra modes for RP, which propagate towards the left edge (positive $\theta$), do not offer an alternative deformation mechanism for the units within the bulk. The only mode available for these is associated with the uniform GH mode, leading to no significant stress localization at the crack tip. However, notable reduction in bending stress is observed along the right edge of RP1-2 (Fig. 5i,j), due to that edge’s FM which locally increase the compliance.
@@
Finally, the stress within the bulk of stretching-dominated lattices (i.e., LUP1) is not significantly affected by the FM introduced at the crack edge. However, at the crack-tip the FM are actuated in a localised manner (Fig. 5k) and these greatly influence the failure, which will be discussed in the following section.
IV.2 Dominant Lattice Failure
@@
Fig. 6a provides the axial stress fraction in the element that fails in the step immediately before fracture. The damage paths, provided in Fig. 6b-g, each showing results for a range of SR, appear to be influenced by the proportion of axial and bending stress within the element. When bending or stretching dominates the fracture, the crack appears to propagate horizontally parallel to the pre-crack edge for symmetric about the mid-plane lattices, i.e. RK at SR = 6, LTK at SR = 2-30, LP1 at SR = 6-30, and RP1 at SR = 20-30. As the stress contributions become similar, the damage path generally deviates from the pre-crack plane as it traverses the lattice.
@@
Figure 6: Dominant failure of lattices. (a) Axial stress fraction in member that fractures preceding the first failure event. (b) The damage process is visualised with a line traversing the domain, connecting separated nodes (fracture locations) within the lattice, indicated by yellow markers. For clarity we do not show the microstructure throughout the domain. The damage process for SR = 3, 6, 10, 30 are plotted with green, purple, black, and blue lines and associated rectangular markers respectively (refer to (a)). Arrows indicate the direction of damage and circular markers indicate a discontinuity in the damage path. Damage paths for: (c) RK; (d) LTK; (e) LP1; (f) RP1; and, (g) LUP1. The simulations for these lattices are included in movie S1.
@@
While the LUP1 global behaviour is stretching-dominated, the fracture of its elements is strongly influenced by the bending (Fig. 6a) associated with the FM localising at the pre-crack edge at large SR (Fig. 7a,d). As such, strong alignment of the damage path with the long-wavelength limit of $\theta_{\mathrm{2}}$ (Fig. 7i) is observed. Simulating other lattices from the equivalence class (-,+,-) at SR = 30, it indicates this is a robust phenomenon: for LUP2 (-0.18,0.04,-0.1) and LUP3 (-0.16,0.04,-0.16) the damage path (Fig. 7g) follows the long-wavelength limit of $\theta_{2}$ and $\theta_{1}$ respectively (Fig. 7d-f,i). This results in the damage evolution following a shallower angle than LUP1. Close to the long-wavelength limit, $\theta_{1}$ propagates further into the bulk for LUP1-3. While the lattice appears to favour propagating the crack in the direction with the smallest $\kappa^{\mathrm{FM}}$ (i.e. LUP3), if $\theta$ is behind the crack tip, it instead favours the modes that propagate forward, ahead of the crack tip (i.e. LUP1 and 2). This phenomenon is also observed for the RDP lattices, provided in Appendix E. For lattices in the ($\pm$$\mp,$$\pm$) equivalence classes, at large SR, there is a strong alignment of the damage path with the FM localised at the crack tip, providing a control parameter to steer the localised fracture of its underlying microstructure. As the SR reduces, the crack tends towards a horizontal path (Fig. 6g), with these FM having a reduced influence on the damage.
@@
Figure 7: Damage path control using FM. Maximum normalised bending stress in the elements cross section $\sigma_{\mathrm{\textbf{Bend}}}=|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$ and visualisation of FM through the magnitude of the total modulation ($|\phi|$) of the wave spectrum for: (a, d) LUP1 (-0.1,0.1,-0.1); (b, e) LUP2 (-0.18,0.04,-0.1); and, (c, f) LUP3 (-0.16,0.04,-0.16). (g) Damage path for LUP1, LUP2, LUP3 at SR = 30. (h) Damage path for lattice with a domain wall separating RDP1 and LUP1. (i) $\theta$ spectrum for LUP1, LUP2 and LUP3. All figures associated with simulations are for SR = 30. The simulations for these lattices are included in movie S3.
V Effect of Discontinuities
@@
V.1 Topological domain walls
@@
Figure 8: Effect of domain walls. Normalised applied stress vs nominal strain for: (a) SR = 3; and, (b) SR = 30. Maximum normalised stress in the elements cross section for D-FM, D-SSS1 and D-SSS2 at specified nominal strains for: (i,iii) SR = 3; and, (ii,iv-viii) SR = 30. Where $\sigma_{\mathrm{M}}=\sigma_{\mathrm{A}}/\sigma_{\mathrm{A}}^{\mathrm{u}}\pm|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$. In (v), (vi), and (viii) yellow stars indicated nodes that are added into the system due to the separation of elements (element fracture) along the domain wall. For clarity, an insert is also included where circular yellow markers indicate these nodes. The simulations for these lattices are included in movie S4.
Pairing lattices from different equivalence classes provides a mechanism to arrest the crack at the domain wall. At SR = 3, LP1 has a higher peak stress than RP1 (Fig. 8a). Once initiated, the crack traverses D-FM until it meets the domain wall (Fig. 8i). The load then increases before the damage continues, resulting in the stiffening behaviour observed in Fig. 8a. In contrast, at SR = 30, RP1 has a higher peak stress, resulting in the damage propagating unstably through D-FM with a sudden reduction in stress (Fig. 8b). This stiffening is also observed for D-SSS1 and D-SSS2 once the damage reaches the SSS domain wall (refer to movie S4).